Find the equations of the spheres which pass through the circle x^2+y^2+z^2=9, 2x+2y-7=0 and the touch the plane x-y+z+3=0
1
Expert's answer
2019-04-09T10:00:18-0400

1. Suppose that the sphere O1 passes through the circle O2 (x2 + y2 + z2 = 9, 2x + 2y - 7 = 0).

Then its equation is:

x2 + y2 + z2 - k(2x + 2y - 7) = 9;

(x - k)2 + (y - k)2 + z2 = 2k2 - 7k + 9,

where k is any real number.

Its center is O1(k, k, 0) and the radius is R1 = "\sqrt{2k^2 - 7k + 9}"


2. Since the sphere touches the plane b (x - y + z + 3 = 0), then the distance O1H from the point O1 to the plane is equal to the radius R1. Point-plane distance is determined by the formula:

Plane P: ax + by + cz + d = 0;

Point O: (x0, y0, z0);

d = |ax0 + by0 + cz0 + d|/"\sqrt{a^2 + b^2 + c^2}"

If a = 1, b = -1, c = 1, d = 3, x0 = k, y0 = k, z0 = 0, then:

d = |1 * k - 1 * k + 1 * 0 + 3|/"\sqrt{1^2 + (-1)^2 + 1^2}" = "3 \over \sqrt3" = "\sqrt3" .

4. From R1 = d:

"\sqrt{2k^2 - 7k + 9}=\sqrt3";

2k2 - 7k + 9 = 3;

2k2 - 7k + 6 = 0;

D = 72 - 4 * 2 * 6 = 49 - 48 = 1;

k = (7 ± 1)/4;

k1 = 1.5; k2 = 2.

The equations of two spheres are:

(x - k)2 + (y - k)2 + z2 = 2k2 - 7k + 9,

1) k = 1.5; (x - 1.5)2 + (y - 1.5)2 + z2 = 3;

2) k = 2; (x - 2)2 + (y - 2)2 + z2 = 3.

Answer: (x - 1.5)2 + (y - 1.5)2 + z2 = 3; (x - 2)2 + (y - 2)2 + z2 = 3.